how do you solve 2y over 3 - 14 over 15 = y over 5
Answer by unlockmath(1688)

Hello,
Let's see if I got this: 2y over 3 - 14 over 15 = y over 5
written as:
(2y/3)-14/15=y/5
IF this is correct then multiply by 15 to get:
10y-14=3y
Subtract 10y to get:
-14=-7y
Divide by -7 and we get:
y=2
Make sense?
RJ
www.math-unlock.com